# DPStyler: Dynamic PromptStyler for Source-Free Domain Generalization

### v1

@@ IV-D Ablation Study
TABLE VI: Ablation studies on DPStyler evaluate Style Generation (SG), Style Remover (SR), and Model Ensemble (ME) across PACS, VLCS, OfficeHome (OH), and DomainNet (DN).
@@
TABLE VII: Ablation experiments on style refresh strategies across PACS, VLCS, OfficeHome (OH), and DomainNet (DN).
Style Generation. We conduct experiments to evaluate DPStyler’s style generation module. As shown in the second row of Table VI, it’s evident that our classifier’s performance has been improved by employing the style generation module alone. This suggests that our style generation module effectively provides more diverse features to the classifier, thereby enhancing its robustness and generalization ability. Furthermore, in order to show that dynamic update styles make the model performs better than fixed styles, we compare the performances of PromptStyle and DPStyler on the four datasets without using style remover and with model ensemble for both. As can be seen from Figure 8, our dynamic style generation module is able to obtain better results than PromptStyler with less training resources. Additionally, we conduct separate tests for using Random only, using StyleMix only, and Random-Mix ($50\%$ Random, $50\%$ StyleMix), as shown in Table VII.
We also explore other structures for the style removal module, as seen in Table VIII. Since the module processes the output vectors of the encoder, the structure should not be too complex. We attempted to use a fully connected layer as the style remover and compared it with the Style-SE Net proposed in this paper. It can be observed that a fully connected layer can also achieve the effect of style removal, but the performance of Style-SE Net is superior.
Head Accuracy (%) $\uparrow$
PACS VLCS OfficeHome DomainNet Avg.
MLP 93.3 82.5 71.9 47.2 73.7
Style-SE Net 93.6 83.5 72.5 48.0 74.4
TABLE VIII: Ablation experiments on different model heads across PACS, VLCS, OfficeHome, and DomainNet. Here ’MLP’ represents a fully connected layer as a style remover, and ’Style-SE Net’ represents our method.
Template Accuracy (%) $\uparrow$
PACS VLCS OH DN Avg.
a [class] in a $S_{*}$ style 93.9 83.2 71.5 47.3 74.0
a $S_{*}$ style of a [class] 93.1 77.0 73.0 47.9 72.8
a photo of a [class] with $S_{*}$ like style 93.7 83.9 70.8 45.7 73.5
Ensemble 93.6 83.5 72.5 48.0 74.4
TABLE IX: The ablation experiments on the Model Ensemble Module using three text templates across PACS, VLCS, OfficeHome (OH), and DomainNet (DN). Ensemble represents the performance of an integrated classifier trained with three different text templates.
Model Ensemble. In Table VI, the fourth and fifth rows illustrate the performance evaluation of an integrated classifier trained using three different text templates. The table demonstrates a significant improvement in overall model performance through model ensemble. Furthermore, Table IX highlights that while integrating classifiers trained with multiple text templates may not yield optimal performance for every dataset, it notably enhances overall performance stability. This enhancement reduces the model’s susceptibility to significant fluctuations resulting from variations in text templates.

### v2

@@ IV-D Ablation Study
TABLE VI: Ablation studies on Style Generation (SG), Style Remover (SR), and Model Ensemble (ME) across PACS, VLCS, OfficeHome (OH), and DomainNet (DN).
@@
Style Generation. We conduct experiments to evaluate DPStyler’s style generation module. As shown in the second row of Table VI, it’s evident that our classifier’s performance has been improved by employing the style generation module alone. This suggests that our style generation module effectively provides more diverse features to the classifier, thereby enhancing its robustness and generalization ability. Furthermore, in order to show that dynamic update styles make the model performs better than fixed styles, we compare the performances of PromptStyler and DPStyler on the four datasets without using style remover and with model ensemble for both. As can be seen from Figure 8, our dynamic style generation module is able to obtain better results than PromptStyler with less training resources. Additionally, we conduct separate tests for using Random only, using StyleMix only, and Random-Mix ($50\%$ Random, $50\%$ StyleMix), as shown in Table IX.
TABLE VII: Ablation experiments on different model heads across PACS, VLCS, OfficeHome, and DomainNet. Here ’MLP’ represents a fully connected layer as a style remover, and ’Style-SE Net’ represents our method.
Head Accuracy (%) $\uparrow$
PACS VLCS OfficeHome DomainNet Avg.
MLP 93.3 82.5 71.9 47.2 73.7
Style-SE Net 93.6 83.5 72.5 48.0 74.4
We also explore other structures for the style removal module, as seen in Table VII. Since the module processes the output vectors of the encoder, the structure should not be too complex. We attempted to use a fully connected layer as the style remover and compared it with the Style-SE Net proposed in this paper. It can be observed that a fully connected layer can also achieve the effect of style removal, but the performance of Style-SE Net is superior.
TABLE VIII: The ablation experiments on the Model Ensemble Module using three text templates across PACS, VLCS, OfficeHome (OH), and DomainNet (DN). Ensemble represents the performance of an integrated classifier trained with three different text templates.
Template Accuracy (%) $\uparrow$
PACS VLCS OH DN Avg.
a [class] in a $S_{*}$ style 93.9 83.2 71.5 47.3 74.0
a $S_{*}$ style of a [class] 93.1 77.0 73.0 47.9 72.8
a photo of a [class] with $S_{*}$ like style 93.7 83.9 70.8 45.7 73.5
Ensemble 93.6 83.5 72.5 48.0 74.4
TABLE IX: Ablation experiments on style refresh strategies across PACS, VLCS, OfficeHome (OH), and DomainNet (DN).
@@
Model Ensemble. In Table VI, the fourth and fifth rows illustrate the performance evaluation of an integrated classifier trained using three different text templates. The table demonstrates a significant improvement in overall model performance through model ensemble. Furthermore, Table VIII highlights that while integrating classifiers trained with multiple text templates may not yield optimal performance for every dataset, it notably enhances overall performance stability. This enhancement reduces the model’s susceptibility to significant fluctuations resulting from variations in text templates.
@@
TABLE XI: The evaluation of training and inference resources on PACS, VLCS and OfficeHome encompasses GPU memory usage, training time, model parameter count, and inference speed. Training Stage 1 represents the training of style word vectors by PromptStyler, and Training Stage 2 denotes the training of the classifier. Training Total represents the cumulative cost of both stages. The symbol ‘*’ denotes reproduced results. ‘ME’ stands for model ensemble, while ‘w/’ and ‘w/o’ respectively indicate whether model ensemble is used.
Method Training Stage 1 Training Stage 2 Training Total Inference
Memory $\downarrow$ Time $\downarrow$ Memory $\downarrow$ Time $\downarrow$ Memory $\downarrow$ Time $\downarrow$ # Params $\downarrow$ FPS $\uparrow$
PACS (7 classes)
ZS-CLIP [30] 0GB 0s 0GB 0s 0GB 0s 102.01M 29.2
PromptStyler* [27] 3.8GB 259s 2.5GB 85s 3.8GB 344s 38.32M 70.0
DPStyler w/o ME 0GB 0s 2.7GB 74s 2.7GB 74s 39.37M 71.2
DPStyler w/ ME 0GB 0s 2.7GB 219s 2.7GB 219s 41.49M 70.6
VLCS (5 classes)
ZS-CLIP [30] 0GB 0s 0GB 0s 0GB 0s 102.01M 30.7
PromptStyler* [27] 3.8GB 245s 2.5GB 69s 3.8GB 314s 38.32M 74.9
DPStyler w/o ME 0GB 0s 2.7GB 70s 2.7GB 70s 39.37M 70.0
DPStyler w/ ME 0GB 0s 2.7GB 203s 2.7GB 203s 41.48M 68.5
OfficeHome (65 classes)
ZS-CLIP [30] 0GB 0s 0GB 0s 0GB 0s 102.01M 10.9
PromptStyler* [27] 4.2GB 679s 2.5GB 346s 4.2GB 1025s 38.91M 69.1
DPStyler w/o ME 0GB 0s 2.7GB 338s 2.7GB 338s 39.43M 70.2
DPStyler w/ ME 0GB 0s 2.7GB 1004s 2.7GB 1004s 41.67M 66.8
IV-E Further Analysis
Non-stylized Domain Shift. When the differences between domains are primarily non-stylized, deeper-level features, we experiment with the DG benchmark on the NICO++ [80] dataset to verify the effectiveness of our method. Some examples of this dataset are given in Fig. 10. The images in the NICO++ dataset are all from the real world, and the differences between domains are different environments. We can see that the same object can be in a picture with grass as the background or in a picture with water as the background in Fig. 10, and that both the grass and the water belong to the non-stylized, deeper-level features. We test the performance of DPStyler on six domains on the DG benckmark  [^2] The DG benchmark for NICO++ can be found here. of this dataset and compare it with the performance of zero-shot CLIP [30]. It can be seen from TABLE XII that DPStyler can effectively improve the classification accuracy.
Figure 10: Example Images of common domains in NICO++. “Autumn”, “Dim”, etc. denote domains, and we need to recognize the category of the objects in the images of each domain
Resource Usage. As shown in Table XI, we provide detailed resource usage comparison between zero-shot CLIP [30], PromptStyler [27] and DPStyler (ours) on PACS, VLCS and OfficeHome, encompassing both training and inference resources. From the table, it’s noticeable that during the training stage without utilizing model ensemble, DPStyler’s training time is only $20$% to $30$% of that required by PromptStyler (as depicted in Table XI’s training total), with a notably significant reduction in GPU memory usage. Remarkably, under this condition, DPStyler achieves an accuracy of $74.0$%, surpassing PromptStyler’s accuracy of $73.3$%. Furthermore, when employing model ensemble, both training time and GPU memory usage remain lower than that of PromptStyler, with consistently superior performance ($74.4\%$ accuracy). Finally, during the testing stage, despite a slight increase in parameter count, our inference speed aligns closely with PromptStyler’s and significantly outperforms ZS-CLIP. Note that there is some randomness in the inference speed considering hardware factors, but in general “DPStyler w/o ME” and “PromptStyler” have similar inference speed.
Style refresh. We conduct an additional ablation study about style refresh, as presented in TABLE XIII. Within this study, we compare our standard DPStyler with a variant where we disable the refresh strategy, meaning the style vectors remain unchanged after their initial setup. With the style remover in place and model ensemble applied, we observe the impact of solely disabling the refresh mechanism on the results. It can be seen that without refreshing styles at the beginning of each training epoch, the model’s performance declines across various datasets. This indicates that refreshing style vectors can generate a more diverse range of styles.
TABLE XII: Comparison with zero-shot CLIP for per-domain top-1 accuracy on NICO++. ZS-CLIP (C) refers to zero-shot CLIP using ‘[class]’ as its text prompt, and ZS-CLIP (PC) indicates zero-shot CLIP using ‘a photo of a [class]’ as its text prompt. Bold indicates the best result.
Method Accuracy (%) $\uparrow$
 Autumn Dim Grass Outdoor Rock Water Avg.
ResNet-50 with pre-trained weights from CLIP
ZS-CLIP (C) [30]  79.9 74.7 84.1 78.9 81.0 68.4 77.8
ZS-CLIP (PC) [30]  81.9 77.2 86.1 80.5 83.0 71.6 80.1
DPStyler 83.6 79.4 87.1 83.7 84.5 76.0 82.4
ViT-B / 16 with pre-trained weights from CLIP
ZS-CLIP (C) [30]  89.7 86.6 90.9 85.7 88.7 80 86.9
ZS-CLIP (PC) [30]  90.3 88.5 91.6 86.8 90.1 82.0 88.2
DPStyler 91.1 88.5 91.4 87.5 90.6 82.4 88.6
ViT-L / 14 with pre-trained weights from CLIP
ZS-CLIP (C) [30]  92.0 89.0 92.2 85.9 90.1 83.7 88.8
ZS-CLIP (PC) [30]  93.0 91.3 92.8 88.2 92.4 85.3 90.5
DPStyler 93.8 91.9 93.0 88.0 93.3 86.6 91.1
TABLE XIII: Ablation experiments on refreshing across PACS, VLCS, OfficeHome (OH), and DomainNet (DN). Bold indicates the best result.
Refresh Strategy Accuracy (%) $\uparrow$
PACS VLCS OH DN Avg.
w/o refresh 93.3 82.8 72.1 47.1 73.8
w refresh 93.6 83.5 72.5 48.0 74.4